# TReMu: Towards Neuro-Symbolic Temporal Reasoning for LLM-Agents with Memory in Multi-Session Dialogues

### v1

@@ Abstract
Temporal reasoning in multi-session dialogues presents a significant challenge which has been under-studied in previous temporal reasoning benchmarks. To bridge this gap, we propose a new evaluation task for temporal reasoning in multi-session dialogues and introduce an approach to construct a new benchmark by augmenting dialogues from LoCoMo and creating multi-choice QAs. Furthermore, we present TReMu, a new framework aimed at enhancing the temporal reasoning capabilities of LLM-agents in this context. Specifically, the framework employs time-aware memorization through timeline summarization, generating retrievable memory by summarizing events in each dialogue session with their inferred dates. Additionally, we integrate neuro-symbolic temporal reasoning, where LLMs generate Python code to perform temporal calculations and select answers. Experimental evaluations on popular LLMs demonstrate that our benchmark is challenging, and the proposed framework significantly improves temporal reasoning performance compared to baseline methods, raising from 29.83 on GPT-4o via standard prompting to 77.67 via our approach and highlighting its effectiveness in addressing temporal reasoning in multi-session dialogues.
@@
Yubin Ge1111thanks: Work done as an intern at AWS, Salvatore Romeo2222, Jason Cai2222, Raphael Shu2222, Monica Sunkara2222, Yassine Benajiba2222, Yi Zhang2222 $1$University of Illinois Urbana Champaign, USA $2$Amazon Web Services yubinge2@illinois.edu {romeosr, cjinglun, sunkaral, zhongzhu, sunkaral, yizhngn}@amazon.com
@@ 1 Introduction
In the context of multi-session dialogues, temporal reasoning is both critical and challenging for LLM-agents. As the number of dialogue sessions increases, storing and retrieving relevant information efficiently becomes more difficult Maharana et al. (2024). LLMs often face challenges in managing large, long-term dialogues, such as failing to retrieve specific temporal details from long history and dialogues exceed the input limit of LLMs. Additionally, research has shown that LLMs overlook important contextual information from long dialogue histories due to the accumulation of irrelevant historical data, referred to as "historical noise" Wang et al. (2023). These challenges underscore the need for enhanced temporal reasoning capabilities in LLM-agents for effective handling of multi-session dialogues.
@@
However, most existing temporal reasoning benchmarks cannot be used directly for this study, because they are built on shorter texts, such as stories and Wikipedia articles, that contain clear temporal information Chen et al. (2021); Wang and Zhao (2024); Xiong et al. (2024). Even benchmarks designed for dialogues, like TimeDial Qin et al. (2021) and LoCoMo Maharana et al. (2024), do not explicitly account for the special temporal characteristics in multi-session dialogues, such as relative time and cross-session dependency. For instance, speakers often use relative time expressions instead of specific dates, requiring the model to infer exact event times. Moreover, it is common for speakers to recall past events from previous sessions, creating cross-session dependencies, where events from different sessions involve the same or related entities and reflect changes over time. This further requires LLMs to retain context effectively when reasoning about events across multiple sessions.
@@ 2 Benchmark Construction
Question Type Count ##\## of Options ##\## of Events Event Type
@@
Table 2: Dataset statistics and details of the constructed benchmark.##\## of Options refers to the number of options for each temporal question. ##\## of Events refers to the number of selected events to create each temporal question. Event Type specifies the type of temporal events chosen for question creation, where (+ relative time) indicates that relative time was an additional consideration in event selection. Note that our benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions.
@@ 2.1 Benchmark Design
We propose augmenting an existing multi-turn dialogue dataset to create a benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues. After a thorough examination, we selected LoCoMo Maharana et al. (2024), which comprises dialogues averaging 600 turns and 16,000 tokens across up to 32 sessions. In comparison to existing multi-session dialogue datasets, LoCoMo features the longest dialogues and the most sessions, thus presenting a greater challenge. The detailed statistics of LoCoMo and its comparison with other relevant datasets are shown in Table 1.
@@ 2.2 Construction Pipeline
To construct our benchmark, we follow the design of our benchmark and utilize a systematic step-by-step approach with GPT-4o. The prompt for each step is shown in Appendix Sec.A.
@@
Step 4: Quality Control We observe various noises in generated QAs, such as incorrect inferences about exact times or the selection of events with ambiguous time expressions (e.g., "the other day"). To ensure the benchmark’s quality, we follow recent temporal reasoning benchmarks for LLMs, such as TGQA Xiong et al. (2024), to perform quality control. We manually review each question to verify that it aligns with our design specifications and that the answers are correctly grounded in the dialogue. We also revise well-constructed questions with incorrect answers and remove any unreasonable ones. The final temporal QA benchmark covers time intervals from days to months and its statistics and details are presented in Table 2. Particularly, our final benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions, which are not covered in LoCoMo. We also include examples of QAs for different temporal reasoning types in Appendix Sec. B.
@@ 3.1 Preliminary: Memory-Augmented LLM-Agents
To address the limit of LLMs struggling in retaining information from long input text, recent studies have demonstrated that equipping LLM agents with memory can efficiently support long-turn conversations Lu et al. (2023); Packer et al. (2023); Zhong et al. (2024). Therefore, we base our study on memory-augmented LLM-agents.
The general pipeline of memory-augmented LLM-agents comprises three stages: memorization, retrieval, and response. In the memorization stage, the model summarizes each dialogue session and stores these summaries as memory. During the retrieval stage, the model retrieves the most relevant memory for the current dialogue session. This retrieved memory is then concatenated with the ongoing dialogue to generate the next response.
For temporal reasoning in multi-session dialogues, we can apply this pipeline by summarizing each dialogue session, retrieving relevant memory for each temporal question, and using this information to select the answer. Specifically, we build our framework based on MemoChat Lu et al. (2023), which effectively realizes this three-stage process through prompting and has demonstrated effectiveness in handling long-range dialogues.
@@ 3.2 TReMu
  Python Executor $\mathcal{P}$
@@
     $\mathcal{M}\leftarrow\mathcal{M}\cup m_{i}$
@@ 3.2.1 Time-aware Memorization
As LLMs may struggle with capturing temporal information in dialogues, particularly relative time, we propose a time-aware memorization strategy using timeline summarization Steen and Markert (2019); Rajaby Faghihi et al. (2022); Sojitra et al. (2024). Specifically, for each dialogue session tagged with its timestamp, we prompt the LLM to generate its summary while also summarizing mentioned events occurring on different inferred dates. The prompt is shown in Appendix Sec.C. This allows us to obtain a series of summaries for a single dialogue session, with each summary is linked to a specific date. By distinguishing between when an event occurred and when it was mentioned, this time-aware memorization could improve temporal reasoning and enable more precise event tracking across the dialogue sessions.
@@ 3.2.2 Neuro-symbolic Temporal Reasoning
Allen’s interval algebra Allen (1983) provides a foundation for solving various types of temporal reasoning by identifying event timestamps and performing temporal comparisons or calculations. For example, determining the difference between the timestamps of two events allows us to resolve temporal interval problems. Building on the interval algebras, previous works have developed various neuro-symbolic models to integrate symbolic reasoning with neural networks Garcez and Lamb (2003); Zhou et al. (2021).
Additionally, researchers have explored neuro-symbolic reasoning for LLMs. By providing LLMs with clear instructions about the grammar of the symbolic language and offering a few demonstrations as in-context examples, LLMs can accurately translate problems into different symbolic language, such as first-order logic Han et al. (2022). Once the problem is translated, symbolic solvers can be employed to solve it, and it has been demonstrated to outperform methods like CoT for multi-hop reasoning tasks Pan et al. (2023).
Inspired by the above motivations, we propose leveraging LLMs to translate temporal reasoning questions into Python code, which can then be executed to derive answers. The detailed prompts are shown in Appendix Sec.D. The underlying reasons stem from the strong ability of state-of-the-art LLMs in generating Python code in multiple code generation benchmarks and by the availability of Python libraries, such as datetime and dateutil, which support temporal calculations. Particularly, dateutil provides a function relativedelta supporting relative time calculation, for example we can infer next Friday using TODAY + relativedelta(weekday=FR). Meanwhile, we provide some implemented functions that the model can directly call, such as "weekRange(t)" returns the start date and the end date of the week that t lies in. As LLMs have familiarity with Python, we provide demonstration via in-context learning to guide it in generating Python code for temporal reasoning, based on the given question and retrieved memory. We then execute the output Python code, and the LLM is prompted again to select the correct answer by using the code.
Particularly, we believe our reasoning approach offers an alternative form of CoT. While the original CoT Wei et al. (2022) performs step-by-step reasoning in natural language, our neuro-symbolic approach conducts temporal reasoning by executing generated code line-by-line in a programming language. This neuro-symbolic method retains the core concept of CoT’s step-by-step reasoning while leveraging the precision and additional support provided by Python code and packages, enhancing the accuracy of the reasoning process.
4 Experiments
4.1 Baselines and Metrics
We build our framework using various backbone LLMs: GPT-4o [^1] Specifically, gpt-4o-2024-05-13., GPT-4o-mini [^2] Specifically, gpt-4o-mini-2024-07-18., and GPT-3.5-Turbo [^3] Specifically, gpt-3.5-turbo-0125.. Particularly, for GPT-3.5-Turbo, many of LoCoMo dialogues are longer than its input length, we then follow LoCoMo Maharana et al. (2024) which earlier dialogues are omitted.
Particularly, in our setting of multi-session dialogues, where conversations may exceed the input limits of LLMs, we also consider the memory mechanism as a critical component of baselines in order to feed complete dialogue information. Therefore, following recent research on temporal reasoning with LLMs Wang and Zhao (2024); Xiong et al. (2024), we compare our approach with the following baseline methods, including direct prompting and LLM-agents with memory:
@@ 4.1 Baselines and Metrics
For evaluation metrics, we primarily use accuracy to assess the overall performance of temporal reasoning. In addition, for unanswerable questions, we calculate precision, recall, and the F1 score to specifically measure performance on this subset of questions. Specifically, precision is computed as the accuracy of questions the model predicts as "unanswerable," while recall is determined by the accuracy of questions where the ground truth answer is "unanswerable."
@@ 4.2 Experimental Results
The results are shown in Tables 3, 4, and 5 for GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo, respectively. On the recent benchmark TRAM Wang and Zhao (2024), existing LLMs have shown strong performance via direct prompting. For instance, on TRAM, GPT-4 achieves an accuracy of 82 with CoT, and GPT-3.5 reaches 71.40 using CoT. In contrast, our benchmark poses a greater challenge, as GPT-4o achieves only 61.67 with CoT and 29.83 with SP, while GPT-3.5’s performance is significantly lower, with 25.83 using CoT and 23.83 with SP. These gaps are likely due to the challenging context of multi-session dialogues and the temporal characteristics in multi-session dialogues, which are not explicitly covered by previous benchmarks.
Besides, our proposed framework outperforms all baseline methods across all three LLMs in terms of both accuracy and F1 score, with a notable increase in accuracy from 29.83 with SP to 77.67 with our framework using GPT-4o. This demonstrates the effectiveness of our approach in enhancing temporal reasoning for multi-session dialogues. A comparison between MemoChat + CoT and Timeline + CoT reveals that the form of memory matters. Time-aware memorization brings higher accuracy due to timeline summarization, which requires models to infer temporal information during memorization, particularly regarding relative time. Notably, incorporating the memory mechanism performs worse than CoT on GPT-4o and GPT-4o-mini. This may be because these models’ input lengths are sufficient to handle LoCoMo dialogues, allowing them to identify relevant temporal information without memory augmentation. In contrast, for GPT-3.5, which has a shorter input limit, adding the memory mechanism generally improves performance by enabling the model to retrieve information from complete dialogues rather than truncated ones.
Additionally, including CoT generally improves the performance, consistent with previous findings Wang and Zhao (2024); Xiong et al. (2024). However, changing from CoT to symbolic reasoning, as implemented in our framework, yields further improvements. This can be attributed to the models’ proficiency in generating Python code without losing the advantage of step-by-step reasoning. When models accurately translate questions into Python, the execution of the code avoids errors, offering an advantage over purely neural reasoning. These results align with recent research that combines LLMs with symbolic reasoners for various reasoning tasks Olausson et al. (2023); Pan et al. (2023).

### v2

@@ Abstract
Temporal reasoning in multi-session dialogues presents a significant challenge which has been under-studied in previous temporal reasoning benchmarks. To bridge this gap, we propose a new evaluation task for temporal reasoning in multi-session dialogues and introduce an approach to construct a new benchmark by augmenting dialogues from LoCoMo and creating multi-choice QAs. Furthermore, we present TReMu, a new framework aimed at enhancing the temporal reasoning capabilities of LLM-agents in this context. Specifically, the framework employs time-aware memorization through timeline summarization, generating retrievable memory by summarizing events in each dialogue session with their inferred dates. Additionally, we integrate neuro-symbolic temporal reasoning, where LLMs generate Python code to perform temporal calculations and select answers. Experimental evaluations on popular LLMs demonstrate that our benchmark is challenging, and the proposed framework significantly improves temporal reasoning performance compared to baseline methods, raising from 29.83 on GPT-4o via standard prompting to 77.67 via our approach and highlighting its effectiveness in addressing temporal reasoning in multi-session dialogues.N
@@
Yubin Ge1,21,2thanks: Work done as an intern at AWS, Salvatore Romeo22, Jason Cai22, Raphael Shu22, Yassine Benajiba22thanks: Work done while at AWS, now at Oracle, Monica Sunkara22, Yi Zhang22 11University of Illinois Urbana Champaign, USA 22Amazon Web Services {yubinge, romeosr, cjinglun, zhongzhu, sunkaral, yizhngn}@amazon.com
@@ 1 Introduction
Temporal reasoning, a fundamental cognitive ability, plays a crucial role in human perception Hoerl and McCormack (2019). It encompasses not only basic concepts such as ordering and duration but also more complex functions like task planning and causal relation discovery Vila (1994). Besides, large language models (LLMs) have brought significant advancements to various tasks. These models typically undergo pre-training on extensive text corpora, endowing them with unprecedented accuracy in predicting the next token given some context Ge et al. (2023a). Despite the remarkable advancements in LLMs and their emerging reasoning capabilities Huang and Chang (2023), previous studies have shown that their performance in temporal reasoning remains suboptimal Chu et al. (2024); Qiu et al. (2024).
In the context of multi-session dialogues, temporal reasoning is both critical and challenging for LLM-agents. As dialogue sessions proceed, storing and retrieving relevant information efficiently becomes more difficult Maharana et al. (2024), such as failing to retrieve specific temporal details from long history and dialogues exceed the input limit of LLMs. Additionally, research has shown that LLMs overlook important contextual information from long dialogue histories due to the accumulation of irrelevant historical data, referred to as "historical noise" Wang et al. (2023). These challenges underscore the need for enhanced temporal reasoning capabilities in LLM-agents for effective handling of multi-session dialogues.
@@
However, most existing temporal reasoning benchmarks cannot be used directly for this study, because they are usually built on shorter texts, such as stories and Wikipedia articles, that contain clear temporal information Chen et al. (2021); Wang and Zhao (2024); Xiong et al. (2024). Even benchmarks designed for dialogues, like TimeDial Qin et al. (2021) and LoCoMo Maharana et al. (2024), do not explicitly consider the special temporal characteristics in multi-session dialogues, such as relative time and cross-session dependency. For instance, speakers often use relative time expressions instead of specific dates, requiring the model to infer exact event times. Moreover, it is common for speakers to recall past events from previous sessions, creating cross-session dependencies, where events from different sessions involve the same or related entities and reflect changes over time. This further requires LLMs to retain context effectively when reasoning about events across multiple sessions.
@@ 2 Benchmark Construction
Question Type Count #\# of Options #\# of Events Event Type
@@
Table 2: Dataset statistics and details of the constructed benchmark.#\# of Options refers to the number of options for each temporal question. #\# of Events refers to the number of selected events to create each temporal question. Event Type specifies the type of temporal events chosen for question creation, where (+ relative time) indicates that relative time was an additional consideration in event selection. Note that our benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions.
@@ 2.1 Benchmark Design
We propose augmenting an existing dataset to create a benchmark for evaluating LLM-agents’ temporal reasoning in multi-session dialogues. After a thorough examination, we selected LoCoMo Maharana et al. (2024), which comprises dialogues averaging 600 turns and 16,000 tokens across up to 32 sessions. In comparison to existing multi-session dialogue datasets, LoCoMo features the longest dialogues and the most sessions (as shown in Table 1), thus presenting a greater challenge.
@@ 2.2 Construction Pipeline
To construct our benchmark, we follow the design of our benchmark and utilize a systematic step-by-step approach with GPT-4o. The prompt for each step is shown in Appendix $\S$ A.
@@
Step 4: Quality Control We observe various noises in generated QAs, such as incorrect inferences about exact times. To ensure the benchmark’s quality, we follow recent temporal reasoning benchmarks for LLMs, such as TGQA Xiong et al. (2024), to perform quality control. We manually review each question to verify that it aligns with our design specifications and that the answers are correctly grounded in the dialogue. We also revise well-constructed questions with incorrect answers and remove any unreasonable ones. The final temporal QA benchmark covers time intervals from days to months and its statistics and details are presented in Table 2. Particularly, our final benchmark not only contains more temporal QAs than LoCoMo, but also include unanswerable questions, which are not covered in LoCoMo. We also include examples of QAs for different temporal reasoning types in Appendix $\S$. B.
@@ 3.1 Preliminary: Memory-Augmented LLM-Agents
To address the limit of LLMs struggling in retaining information from long input text, recent studies turn to equip LLM agents with memory to support long-turn conversations Lu et al. (2023); Packer et al. (2023); Zhong et al. (2024). Therefore, we base our study on memory-augmented LLM-agents.
The general pipeline of memory-augmented LLM-agents comprises three stages: memorization, retrieval, and response. In the memorization stage, the model summarizes each dialogue session and stores these summaries as memory. During the retrieval stage, the model retrieves the most relevant memory for the current dialogue session. This retrieved memory is then concatenated with the ongoing dialogue to generate the next response. Specifically, we build our framework based on MemoChat Lu et al. (2023), which realizes this three-stage process through prompting and has demonstrated effectiveness in handling long-range dialogues.
@@ 3.2 TReMu
 Symbolic Solver $\mathcal{P}$
@@
  $\mathcal{M}\leftarrow f_{org}(\mathcal{M},m_{i})$
@@ 3.2.1 Time-aware Memorization
Our time-aware memorization builds on timeline summarization Steen and Markert (2019); Rajaby Faghihi et al. (2022); Sojitra et al. (2024) and it consists of two steps: Temporal Memory Writing and Memory Organization. During Temporal Memory Writing (prompt in Appendix $\S$.C), we instruct LLM agents to generate memory pieces while also extracting and associating mentioned events with inferred dates. Unlike prior approaches that summarize entire sessions holistically, our method produces fine-grained memory pieces linked to specific inferred time markers. As shown in Tables 3 and 4, our memorization outputs memory pieces corresponding to events with inferred time steps that facilitates to mitigate temporal ambiguity. For example, the highlighted texts show that Michelle cooked a meal and later referenced cooking it at different dates. This enables finer temporal granularity, effectively distinguishing events based on inferred time intervals.
Topic Summary
Catching Up Daniel and Michelle catch up on new events in their lives including new jobs, hobbies, and activities.
Hobbies and Daily Rituals Daniel and Michelle discuss their hobbies and rituals like running, ballet, playing guitar, meditation, and cooking.
Cooking and Celebrations Both talk about their cooking experiences and celebrate Daniel’s promotion.
Books and Recommendations Michelle and Daniel discuss books they’ve read and recommend some to each other.
Personal Items with Sentimental Value Michelle and Daniel talk about cameras and a vintage motorcycle with sentimental value.
Table 3: Output memory from MemoChat based on one dialogue session in LoCoMo.
Time Summary
01/28/2020 Daniel and Michelle share updates on their lives… including Michelle starting her Masters in Psychology, Daniel starting a new job where he learns to code and problem-solve, and Michelle’s hobby of ballet, meditation, and journaling… Michelle mentions she made an Italian meal last Saturday and Daniel made salsa for a taco night. Daniel also shared receiving a promotion …
01/27/2020 Daniel received a promotion and celebrated with a dinner at his favorite spot.
01/25/2020 Michelle cooked a delicious Italian meal for her friends, including pasta, garlic bread, and tiramisu.
01/24/2020 Daniel made a huge batch of salsa and hosted a taco night with friends.
01/20/2020 Michelle started her Masters in Psychology.
Table 4: Output memory from Time-aware Memorization based on the same dialogue session in LoCoMo.
Then, we perform Memory Organization on the output memory pieces to maintain long-term memory. We structure memory in a timeline format, grouping events that occur simultaneously and indexing them based on inferred timesteps. This approach enhances the distinction between an event’s occurrence and its mention, reducing temporal ambiguity and improving time-based retrieval. These enhancements mark a significant difference from traditional memorization approaches, supporting efficiency in temporal reasoning.
@@ 3.2.2 Neuro-symbolic Temporal Reasoning
Inspired by the recent progress in neuro-symbolic reasoning for LLMs Han et al. (2022); Pan et al. (2023), we propose leveraging LLMs to translate temporal reasoning questions into Python code as intermediate rationales, which is executed as the reasoning process to derive answers (prompts shown in Appendix $\S$.D). We tried different symbolic languages and finally chose Python because SOTA LLMs are better at generating Python code and there exist Python libraries that support temporal calculations, like datetime and dateutil. Particularly, dateutil provides a function relativedelta supporting relative time calculation, for example we can infer next Friday using TODAY + relativedelta(weekday=FR). Meanwhile, we provide implemented functions to be directly called, such as "weekRange(t)" returns the start date and the end date of the week that t lies in. Different from other works in temporal reasoning based on code executionLi et al. (2023), we enable our LLM agents with function-calling abilities, ensuring correctness and expanding the range of temporal reasoning tasks beyond simple precedence relations.
We provide demonstration via in-context learning to generate Python code with function calling, given then question and retrieved memory. After the generated code is executed, the output and code serve as intermediate rationales, and the LLM is prompted again to give the answer. Particularly, our reasoning approach offers an alternative form of CoT. While the original CoT Wei et al. (2022) performs step-by-step reasoning in natural language, our neuro-symbolic approach conducts temporal reasoning by executing generated code line-by-line in a programming language. This neuro-symbolic method retains the core concept of CoT’s step-by-step reasoning while leveraging the precision and additional support provided by Python code and packages. However, prior works Li et al. (2023) rely solely on solver outputs without providing intermediate justifications.
@@
4 Experiments
4.1 Experimental Setup
Models. We build our framework using various black-box LLMs: GPT-4o [^1] Specifically, gpt-4o-2024-05-13., GPT-4o-mini [^2] Specifically, gpt-4o-mini-2024-07-18., and GPT-3.5-Turbo [^3] Specifically, gpt-3.5-turbo-0125.. Particularly, for GPT-3.5-Turbo, many of LoCoMo dialogues are longer than its input length, we then follow LoCoMo Maharana et al. (2024) which earlier dialogues are omitted.
Particularly, we have also tried different open-source LLMs but most of them cannot handle the long dialogue inputs from LoCoMo, for example only about 10% of dialogues can be fed into Llama-3-70B. And even for those shorter dialogues that can be fed into Llama-3-70B, we notice that the model gets lost and fails to follow instructions, even failing to generate in the desired format. Therefore, we leave the adaptation to LLMs as future work.
Baselines. Since in our setting of multi-session dialogues where the conversations exceed the input limits of LLMs, we consider the memory mechanism as a critical component of baselines in order to feed complete dialogue information. Therefore, we include the following baselines for comparison:
@@ 4.1 Experimental Setup
To better understand the effectiveness of each component in our framework, we evaluate the following variants as baselines for the ablation study:
@@
Comparing MemoChat + CoT and Timeline + CoT allows us to assess the impact of replacing the standard memory mechanism in LLM agents with our time-aware memorization. Additionally, comparing Timeline + CoT with TReMu highlights the effect of replacing CoT with our neuro-symbolic reasoning approach.
Evaluation Metrics. We primarily use accuracy to assess the overall performance of temporal reasoning. In addition, for unanswerable questions, we calculate precision, recall, and the F1 score to specifically measure performance on this subset of questions. Specifically, precision is computed as the accuracy of questions the model predicts as "unanswerable," while recall is determined by the accuracy of questions where the ground truth answer is "unanswerable."
@@ 4.2 Experimental Results
The results are shown in Tables 5, 6, and 7 for GPT-4o, GPT-4o-mini, and GPT-3.5-Turbo, respectively. On the recent TRAM benchmark Wang and Zhao (2024), existing LLMs demonstrate strong performance with direct prompting. For instance, GPT-4 achieves an accuracy of 82 using CoT, while GPT-3.5 attains 71.40. In contrast, our benchmark is significantly more challenging. GPT-4o achieves only 61.67 with CoT and 29.83 with SP, whereas GPT-3.5 performs even worse, scoring 25.83 with CoT and 23.83 with SP. These performance gaps likely stem from the complexity of multi-session dialogues and their temporal dependencies, which are not explicitly addressed in previous benchmarks.
Our framework outperforms all baseline methods across all three LLMs in terms of both accuracy and F1 scores, with a notable increase in accuracy from 29.83 with SP to 77.67 with our framework using GPT-4o. This demonstrates the effectiveness of our approach in enhancing temporal reasoning for multi-session dialogues. However, incorporating a memory mechanism performs worse than CoT for GPT-4o and GPT-4o-mini. This may be because these models have sufficient input lengths to process LoCoMo dialogues, enabling them to identify relevant temporal information without additional memory augmentation. In contrast, for GPT-3.5, which has a shorter input limit, the memory mechanism generally improves performance by allowing the model to retrieve information from memory rather than truncated dialogues.
Furthermore, we find that incorporating CoT generally improves performance, aligning with previous findings Wang and Zhao (2024); Xiong et al. (2024). In particular, CoT encourages models to search for relevant information within dialogues. However, due to the dialogues’ length, the models sometimes generate responses like "I cannot find the mention of …," which hinders their temporal reasoning capabilities. This further underscores the necessity of a memory mechanism to support long dialogue settings.
@@
4.3 Ablation Study
From Tables 5, 6, and 7, the comparison between MemoChat + CoT and Timeline + CoT highlights the importance of memory representation. Time-aware memorization improves accuracy by instructing models to infer temporal information—particularly relative time—during memorization and mitigate temporal ambiguity. Furthermore, replacing CoT with symbolic reasoning, as seen in the comparison between Timeline + CoT and TReMu, leads to additional performance gains. This improvement stems from the models’ ability to generate Python code while retaining the benefits of step-by-step reasoning. This aligns with recent research integrating LLMs with symbolic reasoners for various reasoning tasks Olausson et al. (2023); Pan et al. (2023).